OpenStudy (anonymous):

let y=(e^x+e^x)/2 c. Find an equation of the line tangent to the curve at x = 1. d. Find an equation of the line normal to the curve at x = 1. e. Find any points where the tangent line is horizontal.

OpenStudy (anonymous):

y=(e^x+e^x)/2=e^x For the tangent line you need d/dx e^x=?

OpenStudy (anonymous):

finding dy/dx and d^2y/dx^2 was the first part in case you need it. dy/dx=e^x d^2y/dx^2=e^x

OpenStudy (anonymous):

Correct. What do you need to find an equation for the tangent line at x=1?? (2 data)

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

A point on the line and the slope of the line. Can you find these?

OpenStudy (anonymous):

Are you trying at least?

OpenStudy (anonymous):

well its just a line right?

OpenStudy (anonymous):

oh wait its not a line its a curve

OpenStudy (anonymous):

Find an equation of the line tangent to the curve at x = 1. This is what we try to find. For any straight line you can easily get an equation if you know a point on the line and the slope of the line. This is a special line as it is tangent to the curve.

OpenStudy (anonymous):

at x=1, the curve is at around 2.5?

OpenStudy (anonymous):

If you do not know how e^x looks like, just type into Google: graph e^x Now back to this problem. So e^x and the tangent to the curve at x=1 have some special properties. They have the same slope at that point. (x=1) And they take the same value there

OpenStudy (anonymous):

at x=1 the value is e^1=e You can leave it like that ( I prefer it that way), or \[e \approx 2.71828\]

OpenStudy (anonymous):

So we have found a point!! At x=1 y=e (1,e) Can you find the slope now?

OpenStudy (anonymous):

the slope would be e too right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So what is the equation of the tangent line?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] This is the formula you need here

OpenStudy (anonymous):

ok would it be ex?

OpenStudy (anonymous):

Yes, y=ex

OpenStudy (anonymous):

so that's the answer to the first part? y=ex?

OpenStudy (anonymous):

OK can you see why the solution of d, is y=-ex ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

wait I might be wrong here :)

OpenStudy (anonymous):

Lets do this step by step, first do you know what a normal line is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK so we know that it goes through the same point (1, e) But it wont have slope m=e, but ?

OpenStudy (mathmale):

In other words, Lamp, a tangent line and a normal line to a curve at a certain point (1,e) have slopes that are related to one another. Do you recall what that relationship is?

OpenStudy (anonymous):

no but I think the slope is e^x?

OpenStudy (mathmale):

Lamp, Andras has already helped you to find the slope of the tangent line to the graph of y=e^x. It's e^x evaluated at x=1. Now you're asked to find the slope of the normal line to the graph at the same point. Again, I ask you how the slopes of perpendicular lines are related.

OpenStudy (anonymous):

No, If you need to find the slope of a line perpendicular to another, you need to take the negative reciprocal. So in this case the m will change to -1/e

OpenStudy (mathmale):

If I were in your shoes and I didn't remember, I'd do a quick Internet search for "slopes of perpendicular lines." Andras is again right on track; the key words here are "negative reciprocal."

OpenStudy (anonymous):

@lamp12345 you are very attractive just sayin

OpenStudy (anonymous):

ok with that slope I got -x/e+e+1/e

OpenStudy (anonymous):

Nice, you know maths :D

OpenStudy (anonymous):

yay! is the last part similar the the first 2?

OpenStudy (anonymous):

Last question is a bit different, any guess?

OpenStudy (anonymous):

First try to find the slope of the line that is horizontal

OpenStudy (anonymous):

the normal line is horizontal, so would I find the slope of that?

OpenStudy (mathmale):

Or, in other words, what's the slope of a horizontal line?

OpenStudy (anonymous):

ok I got a slope of 1/3

OpenStudy (anonymous):

I mean -1/3 @Andras

OpenStudy (anonymous):

No that is wrong  Is this horizontal?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

m has to be 0, if you want a horizontal line.

OpenStudy (anonymous):

ok makes sense

OpenStudy (mathmale):

Lamp, please look up the words "horizontal" and "vertical". It's essential that you know the meaning of "horizontal" before y ou try to solve a problem involving a horizontal line. Could you answer these questions: What is the slope of a horizontal line? What is the slope, if any, of a vertical line?

OpenStudy (anonymous):

tangent line is horizontal This tells us that d/dx =0 In this case e^x=0 Can you solve that?

OpenStudy (anonymous):

x would be 0 right?

OpenStudy (anonymous):

Lamp I have to go now so I finish this quick. The above equation e^x=0 has no solutions!!! This means that there is no horizontal tangent line to e^x If you look at the graph of e^x you can believe this easily.

OpenStudy (anonymous):

any number to the power of 0=1 like e^0=1

OpenStudy (anonymous):

ok so there would be no solution.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

@Andras does that mean there's no points and it doesn't exist?

OpenStudy (mathmale):

No. Andras was correct in stating that YOUR GRAPH HAS NO HORIZONTAL TANGENT LINE. Why? Because the equation e^x=0 has no solution.

OpenStudy (mathmale):